OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 15x − 12 g(x) = 3x −2 0.11 −1 −27 0.33 0 −12 1 2 Complete the chart and determine the solution of the equation f(x) = g(x).

OpenStudy (lynfran):

equate it and solve 15x-12=3x 15x-3x-12=0...

OpenStudy (lynfran):

12x-12=0 12x=12 x=12/12 x=?

OpenStudy (lynfran):

@Naenae1999 what is ur answer?

OpenStudy (anonymous):

1?

OpenStudy (lynfran):

correct